OpenStudy (anonymous):

Determine which ordered pair is a solution of y=8x. A. (0,8) B. (-1,8) C. (1.5,10) D. (2,16)

OpenStudy (carlyleukhardt):

A i think

OpenStudy (anonymous):

That's what I was thinking too. Can you explain if possible?

OpenStudy (carlyleukhardt):

Ummm Idk how to put it in words xD

OpenStudy (utterly_confuzzled):

It's A because if you plug the ordered pairs into the equation, that is the only ordered psi that works

OpenStudy (anonymous):

ok i'll tell u if ur answer was correct or not once i submit the test.

OpenStudy (carlyleukhardt):

utterly is right

OpenStudy (utterly_confuzzled):

Scratch that, it's D. I read it wrong the first time

OpenStudy (carlyleukhardt):

wait, it is?

OpenStudy (anonymous):

im utterly confuzzled lol haha

OpenStudy (carlyleukhardt):

xD. im pretty sure its A tho, A just sounds right then the others.

OpenStudy (anonymous):

Im going to get some people to see if they can help.

OpenStudy (carlyleukhardt):

>.< oke

OpenStudy (anonymous):

@james1769 @YK.PAPI @Liam_Monge

OpenStudy (carlyleukhardt):

@Kitten_is_back

OpenStudy (carlyleukhardt):

@misty1212

OpenStudy (james1769):

wat u need

OpenStudy (anonymous):

wats up @Jazzy17999

OpenStudy (anonymous):

lol hey guys could you help with this question?

OpenStudy (utterly_confuzzled):

Y=8x (2,16) 16=8(2) 16=16 (0,8) 8=8(0) 8=0 that wouldn't work because you're saying zero is equal to 8

OpenStudy (anonymous):

D

OpenStudy (utterly_confuzzled):

So whenever you plug in and work it out like I just did, the only one that works is D

OpenStudy (anonymous):

i got d

OpenStudy (james1769):

^

OpenStudy (anonymous):

thnx Guys I appreciate it. @YK.PAPI @Liam_Monge and also @james1769

OpenStudy (james1769):

no problem @jassy17999

OpenStudy (anonymous):

Mhm anytime